OpenStudy (anonymous):

Anyone experiencing extreme lag right now. A ladder 10 ft. long rests against a vertical wall. If the bottom of the ladder slides away from the wall at 1 ft/s. How fast is the angle between the ladder and ground changing when the bottom of the ladder is 6ft from the wall? Any takers?

OpenStudy (anonymous):

Needed to add some math

OpenStudy (anonymous):

\[\int\limits x^2 dx\]

OpenStudy (anonymous):

YES. That and no assistance.

OpenStudy (anonymous):

Hard to assist with the lag :S

OpenStudy (anonymous):

@Luigi0210 HALP

OpenStudy (anonymous):

hopefully this will help

OpenStudy (luigi0210):

It's obviously 2x ;)

OpenStudy (anonymous):

Luigi, I have a problem for you. A ladder 10 ft. long rests against a vertical wall. If the bottom of the ladder slides away from the wall at 1 ft/s. How fast is the angle between the ladder and ground changing when the bottom of the ladder is 6ft from the wall? Show me your calculus skills.

OpenStudy (luigi0210):

*shanks self*

OpenStudy (anonymous):

AND both of you guys are wrong, integrating x^2 dx = x^3/3 + C

OpenStudy (anonymous):

oh... sorry I guess

OpenStudy (anonymous):

Haha no worries, least you tried :)

OpenStudy (anonymous):

 @Luigi0210 Here's a drawing for you. :)

OpenStudy (luigi0210):

I'm only 12, don't make me math ;-;

OpenStudy (anonymous):

Of coarse I'd put you on the spot after your exam is done.